OpenStudy (giitter):

COWIES HELP ME PLEHS! MEDAL+FAN! <3 @awolflover1

OpenStudy (mndean):

what's your question?

OpenStudy (giitter):

I'll get it ^_^

jhonyy9 (jhonyy9):

can i help you ?

OpenStudy (giitter):

ILL TAKE ANYBODY'S HELP!

OpenStudy (giitter):

@jhonyy9

OpenStudy (giitter):

Ill send you the things up close

jhonyy9 (jhonyy9):

what is your question ? for example in case of the last one there is a circle divided in ten parts so this can you writing in this form of division 10/10 = 1 integer yes ? so the second image has 7 white sectors from these 10 and 3 black sectors from these 10 so this will be wrote in case of the white sectors that is a 7/10 part and the black sector case will be wrote 3/10 part do you understand it ?

jhonyy9 (jhonyy9):

or in case of the 4th image there are 10 column from what there are 2 green and 8 white yes ? so how can you rewriting this using division ?

jhonyy9 (jhonyy9):

the colored part you can writing in the 2/10 division and the white part you can rewriting in the 8/10 division do you understand it now ? bc. the integer 10 column note 10/10 = 1 so and 2/10 +8/10 = (2+8)/10 = 10/10 = 1 hope this is understandably now

OpenStudy (giitter):

Thank you ^_^

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye bye